OpenStudy (anonymous):

all edges of a cube are expanding at a rate of 3 centimeters per second. How fast is the volume changing when each edge is 5 centimeters

OpenStudy (anonymous):

can anybody help

OpenStudy (anonymous):

dl/dt=3, v=l^3, dv/dt=3l^2*dl/dt=l^2 dv/dt=25

OpenStudy (anonymous):

Did you understood what I did there?

OpenStudy (anonymous):

no wats da answer

OpenStudy (anonymous):

Ok, do you know derivatives?